OpenStudy (anonymous):

Solve the following equation. Then enter your answer in the space provided using mixed number format. 7(x - 2) = 3(x + 4)

OpenStudy (anonymous):

first distribute the 7 and 3

OpenStudy (anonymous):

7x-14 =3x+12 -3x -3x 4x-14=12 +14 +14 4x=26 x=6.5

OpenStudy (anonymous):

7x-14=3x+12 7x-14+14=3x+12+14 7x=3x+26

OpenStudy (anonymous):

Right, now you need to isolate the x variable on one side of the equal sign. To do this, just just subtract 3x from both sides